A building casts a shadow 232 meters long. At the same time, a pole 4 meters high casts a shadow 16 meters long. What is the height of the building?

Answer by jim_thompson5910(35256)

(Height of building)/(Length of building's shadow) = (Height of pole)/(Length of pole's shadow)

x/232 = 4/16

x/232 = 1/4

x = 232*(1/4)

x = 232/4

x = 58

So the building is 58 meters tall.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------